OpenStudy (anonymous):

activitity title> constant velocity speed exercises Motion a line with constant velocity. 1. if a bus travels north with a constant velocity of 3 m/s what is the displacement after 15 seconds?

OpenStudy (anonymous):

It is traveling 3 meters every second for 15 seconds. \[\frac{3m}{\not s} \times 15\not s = (3\times 15)m = 45 m\]

OpenStudy (anonymous):

i got 1 - 6 question of my hw =)))))

OpenStudy (anonymous):

how can i get that ? no more conversion?

OpenStudy (anonymous):

hello sir what is displacement?

OpenStudy (anonymous):

Displacement is how far you are from where you started

OpenStudy (anonymous):

thank you sir

OpenStudy (anonymous):

do you want to help me? i got 5 question here left =)

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

activitity title> constant velocity speed exercises Motion a line with constant velocity. 1. if a bus travels north with a constant velocity of 3 m/s what is the displacement after 15 seconds? 2.greg is at the school gate, about 25 m on a straight line from the bleachers. The company commander is giving the countdown of 10 seconds for formation at the bleachers.With what velocity should grey run so that he can be in his designated position without being late? 3.A car moves at 60km/hr along a straight highway.how many seconds will it take for the car to travel a distance of 10 km? 4.if it takes 2 hours for a bus to travel 100 km along a straight highway what is its speed? 5.if a van travel 72 km on a straight expressway for 240 minutes,what is its speed? 6.a jeepney or a car at a constant speed of 40km/hr. how many minutes will it take for the jeepney to travel a distance 0f 80 km?

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

exercises : motion along a line with a constant velocity

OpenStudy (anonymous):

So look at number 2 and see if you can figure out how to set it up the way I did the last one.

OpenStudy (anonymous):

How far does greg have to run?

OpenStudy (anonymous):

25m x 10sec ?

OpenStudy (anonymous):

Not quite. The answer you want is a speed right?

OpenStudy (anonymous):

25m

OpenStudy (anonymous):

\[Speed = \frac{Distance}{Time}\] In the previous problem we used this fact to find Distance.. \[\implies Speed \times Time = Distance\]

OpenStudy (anonymous):

But for this problem, we are given a distance and a time, and need to find the speed.

OpenStudy (anonymous):

So to do that we just divide the distance by the time, and we get speed.

OpenStudy (anonymous):

thats the professor teach us im slow about that.. last subject your formula speed = distance and time your correct sir

OpenStudy (anonymous):

So if \[Speed = \frac{Distance}{Time}\] And the time is 10s and the distance is 25m, what speed is needed?

OpenStudy (anonymous):

25m divide 10s right sir?

OpenStudy (anonymous):

Indeed

OpenStudy (anonymous):

2.5 m/s?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

3.A car moves at 60km/hr along a straight highway.how many seconds will it take for the car to travel a distance of 10 km? So again, we will use the fact that \[Speed = \frac{Distance}{Time}\] Where here we are trying to solve for the Time, and we have the speed and the distance.

OpenStudy (anonymous):

Can you solve it?

OpenStudy (anonymous):

sir i got a real question to you sir sorry for disturbing you.the proffesor quiz yesterday and he answer it.....

OpenStudy (anonymous):

can i put it?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

a<<< symbol like a with > in his head what is the a with ">" of a car that goes 20km/ hr to 30 km / hr in 2sec?

OpenStudy (anonymous):

i never thought he use conversion with that ...TT

OpenStudy (anonymous):

I think he means acceleration. \[Acceleration = \frac{\text{Change in speed}}{\text{Change in time}}\]

OpenStudy (anonymous):

So if it goes from 20 to 30 km, what is the change in speed?

OpenStudy (anonymous):

yes sir acceleration sir

OpenStudy (anonymous):

Err 20km/s to 30km/s rather.

OpenStudy (anonymous):

the proffesors answer is 1.80x10 km /hr <<<with small 4 in 10

OpenStudy (anonymous):

he use conversion of 2s x 1km divide 3600s we didnt know he do that

OpenStudy (anonymous):

Because you must have all your units be the same. \[\text{Change in speed} = \frac{30km}{hr} - \frac{20km}{hr} = \frac{10km}{hr}\] \[\text{Change in time} = 2s = 2s \times \frac{1min}{60s} \times \frac{1hr}{60min} = \frac{2}{3600}hr = \frac{1}{1800}hr\] \[Acceleration = \frac{\frac{10km}{hr}}{\frac{1}{1800}hr} = 1.8\times 10^4 \frac{km}{hr^2}\]

OpenStudy (anonymous):

sir so the prof is correct? i thought the prof is lazy i knew it

OpenStudy (anonymous):

your answer is correct sir its written in my notebook

OpenStudy (anonymous):

lets go back to no.3 sir thank you sir

OpenStudy (anonymous):

So, you have the speed, and you have the distance, you have to solve for the time. Can you manage it?

OpenStudy (anonymous):

60 x 10 sir?

OpenStudy (anonymous):

\[Speed = \frac{Distance}{Time}\]\[\implies 60 = \frac{10}{Time}\] Solve for Time.

OpenStudy (anonymous):

600

OpenStudy (anonymous):

No. If you are going 60km each hour it will not take you 600 hours to go 10km

OpenStudy (anonymous):

ops

OpenStudy (anonymous):

60 divide 10 = 6hr = 21600 sec?

OpenStudy (anonymous):

You are going 60 km each hour. Will it take more or less than 1 hour to go 10km?

OpenStudy (anonymous):

less hour 6mins?

OpenStudy (anonymous):

360 sec?

OpenStudy (anonymous):

\[60 = \frac{10}{Time}\]\[\implies \frac{60}{10} = \frac{1}{Time}\]\[\implies 6 = \frac{1}{Time}\]\[\implies Time = \frac{1}{6}hr = \frac{1}{6}{60}s = 10s\]

OpenStudy (anonymous):

thank you sir i feel shy to disturb you maybe you answer 4,5,6 and u can rest and i can rest to TT

OpenStudy (anonymous):

i write your answer to my hw =)

OpenStudy (anonymous):

sir answer of no.4 50km/hr?

OpenStudy (anonymous):

waaaaaaaaaa